OpenStudy (imstuck):

Help please! What is the derivative of y = e^(1+ln(x)) with steps please!!!

OpenStudy (xapproachesinfinity):

Chain rule! this not some fake question! you know how to do it!

OpenStudy (xapproachesinfinity):

i'm suspicious about it lol

OpenStudy (loser66):

agree with @xapproachesinfinity

OpenStudy (imstuck):

Actually I am not getting it correct and I can't figure out why!

OpenStudy (imstuck):

seriously, please help. I can't get it right!

OpenStudy (loser66):

Question: is it not that e^ (a+b) = (e^a )(e^b)?

OpenStudy (xapproachesinfinity):

okay anyways! \(\Large \rm (f(g(x))'=f'(g(x))*g'(x)\) chain rule says take the derivative of the outer with respect to the inner variable g(x) multiplied by the derivative of the inner function with respect to x so \(\huge y=e^{1+lnx}=e*e^{lnx}\) so \(\large y'=e(e^{lnx})'=e*e^{lnx}*(lnx)'=e*e^{lnx}*1/x\)

OpenStudy (xapproachesinfinity):

writing it in another way \(\huge \rm y'=\frac{e^{1+lnx}}{x}\)

OpenStudy (loser66):

One more question: is it not that \(e^{ln(x)}=x\)

OpenStudy (xapproachesinfinity):

hhhh i'm not paying attention lol

OpenStudy (xapproachesinfinity):

so the final product is e

OpenStudy (loser66):

:)

OpenStudy (xapproachesinfinity):

i was focused on derivative i didn't realize it was e^lnx

OpenStudy (imstuck):

ok you guys, I get it now! I didn't remember about the derivative of e^lnx either! That's the whole problem! I don't know who to medal for this. If I medal one of you, the one I medall needs to medal the other guy!!! Thank you both so much!!!!!

OpenStudy (xapproachesinfinity):

writing things neatly \(\huge y'=\frac{e*x}{x}=e\)

OpenStudy (xapproachesinfinity):

eh don't worry about medals I'm fine! the derivative of e^x is just e^x

OpenStudy (imstuck):

Ty for medaling, @Loser66 . what a good sport you are! It's beeen awhile with the e^ something; I forgot! ugh! Thank you both a;gain, so much!!!!

OpenStudy (xapproachesinfinity):

eh this is stupid of me from the start e^lnx=x so the function is f(x)=ex f'(x)=e i'm dumb lol

OpenStudy (loser66):

Thanks for compliment. I can see other's problem only, not mine. I am still struggling with my problem. If you guys can, please, help

OpenStudy (xapproachesinfinity):

welcome^_^

OpenStudy (xapproachesinfinity):

your problem are tough, I forgot how to do those stuff.